OpenStudy (jazzyfa30):

please help Exponential and Logarithmic Equations solve each equation. Give an exact solution. 3^(4x+1)-5=22

hartnn (hartnn):

start by adding 5 on both sides

OpenStudy (jazzyfa30):

3^(4x+1)=27

hartnn (hartnn):

yes, now you can take log on both sides. what do you get ?

hartnn (hartnn):

Alternative way : write 27 as an exponent of 3 and compare the exponents

OpenStudy (jazzyfa30):

can you use draw cus im lost

hartnn (hartnn):

\(\Large 3^{4x+1} = 27 \\ \large \log (3^{4x+1}) = \log 27 \)

OpenStudy (anonymous):

log shmog

hartnn (hartnn):

now use the log property that \(\log a^b = b \log a\)

OpenStudy (anonymous):

\[\large 3^{4x+1}=3^3\iff 4x+1=3\]

hartnn (hartnn):

^ and there's your alternative way

OpenStudy (jazzyfa30):

\[4x+1\log_{3}=\log_{27} \]

OpenStudy (anonymous):

arg

OpenStudy (anonymous):

since \(27=3^3\) you know \(4x+1=3\)

OpenStudy (jazzyfa30):

x=1/2

hartnn (hartnn):

yeah, but did you get how ?

OpenStudy (jazzyfa30):

yes